OpenStudy (anonymous):

Select the correct graph of -x - 2y = 6. graph of line crossing the y axis at 3 and through point parentheses 2 comma 2 close parentheses graph of line crossing the y axis at 3 and through point parentheses 1 comma 1 close parentheses graph of line crossing the y axis at negative 3 and through point parentheses 1 comma negative 5 close parentheses graph of line crossing the y axis at negative 3 and through point parentheses 2 comma negative 4 close parentheses

OpenStudy (anonymous):

@princeharryyy

OpenStudy (anonymous):

@ninamae05 @nincompoop

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@hunterhoff

OpenStudy (anonymous):

im thinkin its a

OpenStudy (anonymous):

-x - 2y = 6 is the same as y= -(1/2)x-3 i just changed it to y=mx+b so from that equation you know that you have slope=-1/2, y-intercept=-3 It can't be a because your y-intercept is negative, and if you substitute (2,2) in the given equation you'll get -6 not 6. so try again :)

OpenStudy (anonymous):

b

OpenStudy (princeharryyy):

if u don't get anything about the ques plug in the values of x & y which ever satisfies will b ur ans @taylor12344

OpenStudy (anonymous):

"graph of line crossing the y axis at 3" NOT RIGHT.. i told you before that y-intercept is NEGATIVE 3, so this cant be the right choice. and p.s. please dont guess:)

OpenStudy (anonymous):

so its c or d

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

hmm.. WHY DO YOU THINK IT IS "C"? *not saying if it is right*

OpenStudy (anonymous):

bc y intersecp[t is -3 and x intersecpt is -1-2

OpenStudy (anonymous):

As part of an Earth Science project, Erin tracked the amount of rain fall for 4 weeks. She started tracking after finding some initial data on the year's rainfall to date: Week Rainfall 1 5.5 2 8 3 10.5 4 13 If the linear function representing the amount of rainfall each week is f(w) = 2.5w + 3, what does the 3 represent in this function? The amount of rainfall in week 0 The week that Erin collected data The amount of total rainfall The amount the rainfall increased each week

OpenStudy (anonymous):

im thinking its a

OpenStudy (anonymous):

how about the point? do you think it is right? You are not done with the first question yet.

OpenStudy (anonymous):

no its d

OpenStudy (anonymous):

WHY ?

OpenStudy (anonymous):

i dk

OpenStudy (anonymous):

*sigh* because (2, -4) satisfy the given equation, y int =-3 and slope is -1/2 next time, please try to understand your topic or try to figure out how to do it because it will help you with your other questions. Don't depend too much on others because they will not last long, you have to depend on yourself. Help yourself :) that's all. D is right.